OpenStudy (anonymous):

Write an equation for the parabola in standard form with (3,2) and (4,0)

OpenStudy (marissalovescats):

Which of these is the Vertex and the Focus?

OpenStudy (anonymous):

vertex 4,0 focus 3,2

OpenStudy (marissalovescats):

Okay give me a sec

OpenStudy (anonymous):

just sent the graph

OpenStudy (anonymous):

apply the definition of parabola.

OpenStudy (dan815):

is standard form ax^2+bx+c or vertex form a(x-k)^2=(y-c)

OpenStudy (amistre64):

since your V and F ... if thats how they are defined ... are not on the same x or y axis, your parabola is skewed

OpenStudy (amistre64):

oh, vertex is 4,0 and a point on it is 3,2

OpenStudy (anonymous):

you have the vertex, but not the focus what you have is another point on the graph you know it looks like \[y=a(x-4)^2\] because you only have one zero at \(x=4\) solve for \(a\) via \[2=a(3-4)^2\]

OpenStudy (anonymous):

the axis of this parabola is not parallel to the x-axis or the y-axis.

OpenStudy (anonymous):

because if \((3,2)\) i on the graph, that means if \(x=3\) then \(y=2\)

OpenStudy (dan815):

amistre is gonna be 100 soon

OpenStudy (amistre64):

only if i keep kicking satellite off the OS ;)

OpenStudy (anonymous):

i think you will find that \(a=2\) and so the equation is for \[y=2(x-4)^2\] or whatever you get when you multiply that out

OpenStudy (amistre64):

sateliite deserves the 100 spot, since he is simply smarter then me

OpenStudy (anonymous):

is that why i keep getting suspended? i thought it was for spam

OpenStudy (anonymous):

not true, just have more time on my hands now that i quit ...

OpenStudy (amistre64):

i have more time on my hands now that i got an office job ...

OpenStudy (anonymous):

vertex-4,0 and focus 3,2. the axis of the parabola is at some inclination to the x axis. it is not like y=ax^2+bx+c

OpenStudy (rsadhvika):

yeah if thats the given data, then it would get complicated i think interesting to see how the equation looks for a skewed parabola...

OpenStudy (rsadhvika):

equation may have sin and cos ?

OpenStudy (amistre64):

if those are the focus and vertex, we could move this so the vertex is at the origin: V = 0,0, F = -1,2, and determine the slope of the line for the tan(a) then rotate it about using X = x cos(a) ... etc

OpenStudy (rsadhvika):



OpenStudy (amistre64):

the distance between them is sqrt(5) so an equivalent parabola with vertex 0,0 and focus 0,sqrt(5) could be constructed then rotated back and shifted again

OpenStudy (marissalovescats):

I'm thinking it's